Grade 12th passPhysical Chemistry

the density of 3M solution of NaCl is 1.25 g/ml. calculate molality of the solution.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer :

For 3M solution NaCl means, 3 moles of NaCl in 1 L of solvent
By, using the density we can find out the mass of the 1 L of solution.
mass of solution= volume x density = 1000 ml (1 L) x 1.25 g/ml = 1250 g
mass of solute = mass of 3 moles of NaCl = 58.5 g/mol x 3 mol = 175.5 g...............(molal mass of NaCl = 58.5 g/mol)
So, mass of solvent = mass of solution - mass of solute = 1250- 175.5 gm = 1074.5 gm = 1.07 Kg
So, molality = moles of solute / mass of solvent = 3 mol/ 1.07 kg = 2.8 mol kg-1

Thank you !